OpenStudy (anonymous):

4!= A. 24 B. 16 C. 1/4 D. 10

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (kirbykirby):

\(4!=4\times 3\times 2\times 1\)

OpenStudy (skullpatrol):

I think it is easier to remember it as 1*2*3*4

OpenStudy (kirbykirby):

Hm we always learnt it in that order hehe. We always saw \(n!=n(n-1)(n-2)\cdots 3\cdot 2\cdot 1\)